# Masked Diffusion Models are Secretly Time-Agnostic Masked Models and Exploit Inaccurate Categorical Sampling

### v3

@@ I.1.1 Comparison of Training Variants
I.1.2 Failed Training Attempts
@@
I.2 Sampling Results
I.2.1 Comparison of High-Order Variants
@@
I.2.2 Impact of Numerical Precision on Our Sampler
@@
In Figure 14, we examine the impact of numerical precision on our first-hitting sampler by varying the floating-point precision in categorical sampling between 32-bit and 64-bit. For the high-order variants, we employ the extrapolation strategy when $N\leq 128$ and the predictor-corrector strategy when $N\geq 256$. In contrast to the observations under MDM’s original sampler, the temperature-lowering effect of the numerical truncation is significantly less influential under our sampler. Notably, the 32-bit second-order first-hitting sampler even results in slightly higher entropy compared to its 64-bit counterpart when $N\geq 512$. This implies that the truncation effect is more pronounced when performing categorical sampling on multiple tokens simultaneously, possibly due to the imbalanced unmasking probabilities. The first-hitting sampler, which ensures that remaining masked tokens are equally unmasked, proves to be more robust to the truncation.
I.2.3 Examples of Generated Text

### v4

@@ Appendix I Old Version of the Introduction
J.1.2 Failed Training Attempts
@@
J.2 Sampling Results
J.2.1 Comparison of High-Order Variants
@@
J.2.2 Impact of Numerical Precision on Our Sampler
@@
In Figure 14, we examine the impact of numerical precision on our first-hitting sampler by varying the floating-point precision in categorical sampling between 32-bit and 64-bit. For the high-order variants, we employ the extrapolation strategy when $N\leq 128$ and the predictor-corrector strategy when $N\geq 256$. In contrast to the observations under MDM’s original sampler, the temperature-lowering effect of the numerical truncation is significantly less influential under our sampler. Notably, the 32-bit second-order first-hitting sampler even results in slightly higher entropy compared to its 64-bit counterpart when $N\geq 512$.
Figure 15: Illustration of prioritized unmasking.
Why the numerical issue is not notable in token-by-token decoding
With our first-hitting sampler, the inference of MDMs becomes a token-by-token decoding process, except that the time variable is additionally handled. In this case, the numerical issue becomes negligible, and 32-bit floating-point precision appears sufficient. We also observe that ARMs, which also adopt a token-by-token sampling strategy, do not suffer from numerical issues under 32-bit precision. This suggests that the numerical problem is a distinctive characteristic of the vanilla sampling process of MDMs (Eqn. (10)) that performs inaccurate categorical sampling simultaneously on all mask positions.
This phenomenon can be explained as follows. As justified theoretically in Section 5.3, the implication of inaccurate categorical sampling includes two aspects: (1) temperature-lowering effect for a single token position and (2) prioritized unmasking for different positions (Figure 15). Both factors reduce the diversity and lower the entropy. However, when altering to token-by-token decoding, all remaining mask tokens have equal probability to be first unmasked, and the diversity decrease becomes less pronounced. This suggests that the prioritized unmasking caused by shifted probabilities at all individual positions is the major factor. This effect accumulates across numerous sampling steps, eventually leading to notable diversity issues, even under 32-bit floating-point precision.
J.3 Anaysis of the Efficiency Gain by Our Sampler
Due to the theoretical equivalence between the FHS and the original MDM sampling procedure, the correctness of the FHS is guaranteed and irrelevant to vocabulary size or sequence length. However, the efficiency gains (measured by inference wall-clock time) depend on several factors.
Let the sequence length be denoted as $L$, vocabulary size as $|V|$ (excluding the mask token), the number of sampling steps (for original MDM sampling) as $N$, the number of function evaluations as $NFE$, and the number of categorical sampling operations as $NCS$. As stated in Section 4, our FHS reduces inference time by minimizing $NCS$. For original MDM sampling with the caching strategy, $NFE\approx N(1-(1-1/N)^{L})$, $NCS=NL|V|$. For the FHS, $NCS=L|V|$. The total time cost can be expressed as $NFE\times t_{1}+NCS\times t_{2}$, where
  • $t_{1}$ is the time for one network call, influenced by model size.
  • $t_{2}$ is the time for categorical sampling, averaged per position and class, which involves two logarithmic operations, and is fixed.
For a fair comparison, we evaluate under the same $NFE$, ensuring similar generation quality. The inference time ratio between original MDM sampling and the FHS is $(NFE\times t_{1}+NL|V|t_{2}):(NFE\times t_{1}+L|V|t_{2})$. Therefore, the FHS yields larger speedups under the following conditions:
  • Smaller Model Size: When the model is smaller, $t_{1}$ decreases, making categorical sampling relatively more expensive compared to network evaluation.
  • Larger $NFE$, $L$, or $|V|$: Higher values for these parameters increase $NCS$ in original MDM sampling, amplifying the speedup provided by the FHS.
Examples:
  • Paper Case: $|V|=50{,}526$, $L=1024$. To match the original MDM sampling at $N=10{,}000$ steps, $NFE\approx 973$. The model size is approximately 600M parameters, with a time ratio of $t_{1}:L|V|t_{2}\approx 1:1.56$. In this case, the inference time ratio is around 17, corresponding to a 20$\times$ speedup. If $N=2048$, the ratio drops to around 5$\times$.
  • DiffSound (Yang et al., 2023): This model is 3$\times$ larger and uses a much smaller vocabulary. Specifically, $L=265$, $|V|=256$, with $t_{1}:L|V|t_{2}\approx 14.6:1$. Here, categorical sampling is much cheaper relative to network evaluations. Using fewer steps, $N=100$, results in $NFE\approx 93$, and the speedup ratio is only around 1.07$\times$.
J.3.1 Examples of Generated Text